Find the cube root of each of the following numbers by prime factorisation method.

 

(iii) 10648

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


(iii) 10648

Solution:

10648 = 2×2×2×11×11×11

By grouping the factors in triplets of equal factors, 10648 = (2×2×2)×(11×11×11)

Here, 10648 can be grouped into triplets of equal factors,

∴ 10648 = 2 ×11 = 22

Hence, 22 is cube root of 10648.





Thanks